# Deep Time Series Models: A Comprehensive Survey and Benchmark

### v1

@@ 6.3.1 Handling Extremely Long Series
Deep time series models have demonstrated remarkable performance across a wide range of downstream tasks, their applicability to longer time series data is often limited by scalability and high computational complexity. In industrial time series analysis, high-frequency sampling results in lengthy historical data, impeding the practical implementation of advanced deep models. Existing methods usually include patching techniques to enable them to handle long sequences, and when the input length becomes longer, the patch length can be increased accordingly to reduce the computational complexity. However, model performance is closely tied to patch length; hence, solely increasing patch size to reduce complexity may compromise capabilities. Therefore, addressing the limitations of deep models in handling longer time series could be a promising topic.
@@ 6.3.2 Utilizing Exogenous Variables
Since variations within the time series are often influenced by external factors, it is crucial to include exogenous variables in the analysis to gain a more comprehensive understanding of these factors. Exogenous variables, which are widely discussed in time series prediction tasks, are included in the model for uniform training in modeling time series data, without requiring separate analysis. However, in practical applications, different from multivariate time series analysis, the main variables and covariates usually occupy different positions. Given the crucial role played by exogenous variables in real-world applications, it is essential to explore a unified framework for modeling the relationships between the endogenous and exogenous variants, which allows for a more comprehensive understanding of interrelations and causality among different variants, leading to better and more reliable model performance, as well as interpretability.
@@ 6.3.3 Processing Heterogeneous Data
As researchers continue to explore this unexplored area in time series analysis, there is potential for significant advances in areas such as finance, healthcare, and environmental monitoring. By improving our ability to model and analyze heterogeneous time series data, we can gain deeper insights into complex systems and make more informed decisions based on predictive analytics. Overall, further research in this area holds great promise for advancing our understanding of temporal data dynamics and enhancing the capabilities of time series modeling in real-world applications.
@@ 7 Conclusion
In this survey, we provide a systematic review of deep models in time series analysis and introduce Time-Series Library (TSLib) as a fair benchmark for deep time series models across various analysis tasks. Compared with previous reviews that focus on a specific analysis task or model architecture, this paper provides a comprehensive survey and overview of existing deep models for time series analysis, ranging from forecasting, classification, imputation, and anomaly detection. We first present a detailed review of the universal modules that are widely used among time series models, including normalization, decomposition, and Fourier analysis. Next, we summarize existing deep time series models from the perspective of backbone architecture. Based on the review of existing literature, we introduce a practical open-source library, Time Series Library (TSLib), which has included representative deep time series models that can be a fair evaluation benchmark in the field of time series analysis. Finally, we discuss future research directions for deep time series models based on the recent development of the AI community and the practical application needs of time series analysis in real-world scenarios.

### v2

@@ 6.3 Practical Applications
6.3.1 Probabilistic Forecasting
Probabilistic forecasting is a pivotal advancement in practical applications of deep time series models, particularly for informed decision-making. Unlike previous widely-studied deterministic forecasting, probabilistic forecasting offers a distribution over potential future outcomes, enabling stakeholders to assess risks, evaluate confidence intervals, and make robust decisions under uncertainty. Most existing probabilistic forecasting models rely on predefined probability distributions and are designed to predict the parameters of the distributions in the future. While these methods have achieved significant success, the dependence on predefined distributions can limit their ability to represent complex and heterogeneous real-world uncertainties. Therefore, accurately modeling and representing such complexities to generate more flexible and robust probabilistic forecasts remains a critical and active area of future research.
@@ 6.3.2 Handling Extremely Long Series
Deep time series models have demonstrated remarkable performance across a wide range of downstream tasks, but their applicability to longer time series data is often limited by scalability and high computational complexity. In industrial time series analysis, high-frequency sampling results in lengthy historical data, impeding the practical implementation of advanced deep models. Existing methods usually include patching techniques to enable them to handle long sequences, and when the input length becomes longer, the patch length can be increased accordingly to reduce the computational complexity. However, model performance is closely tied to patch length; hence, solely increasing patch size to reduce complexity may compromise capabilities. Therefore, addressing the limitations of deep models in handling longer time series could be a promising topic.
@@ 6.3.3 Utilizing Exogenous Variables
Since variations within the time series are often influenced by external factors, it is crucial to include exogenous variables in the analysis to gain a more comprehensive understanding of these factors. Exogenous variables, which are widely discussed in time series prediction tasks, are included in the model for uniform training in modeling time series data, without requiring separate analysis. However, in practical applications, different from multivariate time series analysis, the main variables and covariates usually occupy different positions. Given the crucial role played by exogenous variables in real-world applications, it is essential to explore a unified framework for modeling the relationships between the endogenous and exogenous variables, which allows for a more comprehensive understanding of interrelations and causality among different variants, leading to better and more reliable model performance, as well as interpretability.
@@ 6.3.4 Processing Heterogeneous Data
As researchers continue to explore this relatively uncharted area in time series analysis, there is potential for significant advances in areas such as finance, healthcare, and environmental monitoring. Enhancing models’ ability to analyze heterogeneous time series data can lead to deeper insights into complex systems, enabling more informed decision-making through predictive analytics. Overall, further research in this area holds great promise for advancing our understanding of temporal data dynamics and significantly enhancing time series modeling capabilities in real-world applications.
@@ 7 Conclusion
In this survey, we provide a systematic review of deep models in time series analysis and introduce Time-Series Library (TSLib) as a fair benchmark for deep time series models across various tasks. Compared with previous reviews that focus on a specific analysis task or model architecture, this paper provides a comprehensive survey and overview of existing deep models for time series analysis, ranging from forecasting, classification, imputation, and anomaly detection. By separating basic modules from complicated and diverse deep time series models, we successfully uncover the fundamental design principles for current research, forming an inspiring model-centric perspective. Besides, based on TSLib, which enables flexible and comprehensive model evaluation, we conduct extensive experiments to bring insights for model design and practical usage. With clear summarization, rich experiments, and in-depth discussion, we believe that this survey and TSLib can be instructive and useful for future research on deep time series models.